A single NEC SX-9 vector processor unit (VPU) has a peak arithmetic

performance of 102.4 GFs/s. The memory system supplies a peak memory

bandwidth of 162 GBs/s. The VPU dissipates 40 W, while the memory system

dissipates 55 W. Together, both components dissipate 95 W. Every program

has an arithmetic intensity (AI), measured in flops/byte, that indicates

how much performance is achievable for a given bandwidth. The equation is:

achievable GFs/s = min{AI*BW,peak GFs/s}.

a) What is the minimum AI that achieves peak arithmetic performance on an

SX-9 VPU?

                                                      AI = _____ flops/byte

b) What is the power efficiency of this computer system at peak performance?

                                                         PE = _____ GFs/s/W

c) A special customer buys two memory systems for the one VPU, thus doubling

both the bandwidth and the power of the memory. What is the new minimum AI

that achieves peak arithmetic performance?

                                                      AI = _____ flops/byte

d) What is the power efficiency of the new system at peak performance?

                                                         PE = _____ GFs/s/W

e) Since power efficiency is a good thing, why would a customer spend more

money to get less of it?
Answer #1

Answer:

Given that:

A single NEC SX - 9 vector processor unit (VPU) has arthmatic performance of 102.4 GFs/s.The memory system supplies a peak memory bandwidth of 162 GBs/s

a) Formula Arthemetic Performance = Arthemetic Intensity x Bandwidth

So AI=AP/B

AI=102.4/162

AI=0.63 flops/byte

b) Formula Power Effiency=Arthemtic Performance/voltage

= 102.4/95

PE =1.07 GFs/W

c) AI=AP/B

here Bandwidth is double so B=2B

so AI=102.4/162x2

AI =0.31 Flops/byte

d) Power Effiency=Arthemtic Performance/voltage

Arthemetic Performance = Arthemetic Intensity x Bandwidth

AP=0.31x162x2

AP=100.4

PE=AP/V

PE=100.4/95

PE=1.05 GFs/W

e) IF effiency is more then it will be effective